OpenStudy (anonymous):

Verify the identity. cos 4u = cos22u - sin22u

OpenStudy (anonymous):

4u=2u+2u

OpenStudy (anonymous):

Oh, sorry. The problem is \[\cos4u=\cos^22u-\sin^22u\]

OpenStudy (anonymous):

@Jonask

OpenStudy (anonymous):

use the identity\[\cos 2\theta=\cos^2\theta-\sin ^2\theta\] \[\cos 2(2u)=?\]

OpenStudy (anonymous):

cos 2u=cos^2u-sin^2u then cos 2(2u)= cos^2(2u)-sin^2(2u)

OpenStudy (anonymous):

Oh ok. Then what do I do? Subtract cos^2(2u)?

OpenStudy (anonymous):

@jishan